# Fast and accurate emulation of complex dynamic simulators Junoh Heo

### v1

@@ Abstract
While dynamic simulators, which are computational models that evolve over time and are governed by differential equations, are essential in scientific and engineering applications, their emulation remains challenging due to the unpredictable behavior of complex systems. To address this challenge, this paper introduces a fast and accurate Gaussian Process (GP)-based emulation method for complex dynamic simulators. By integrating linked GPs into the one-step-ahead emulation framework, the proposed algorithm enables exact analytical computations of the posterior mean and variance, eliminating the need for computationally expensive Monte Carlo approximations. This significantly reduces computation time while maintaining or improving predictive accuracy. Furthermore, the method extends naturally to systems with forcing inputs by incorporating them as additional variables within the GP framework. Numerical experiments on the Lotka-Volterra model and the Lorenz system demonstrate the efficiency and computational advantages of the proposed approach. An R package, dynemu, implementing the one-step-ahead emulation approach, is available on CRAN.
@@
Junoh Heo
@@ 1 Introduction
Dynamic systems often have forcing inputs, which are external or time-varying variables that influence the system’s evolution. Examples include external forces in mechanical systems, atmospheric conditions in climate models, or control inputs in engineering applications. These inputs are typically included as additional input variables in the emulator and are usually assumed to be known. Their presence modifies the system dynamics, necessitating an adaptation of the emulation approach to capture their effects properly.
@@
GPs (Rasmussen and Williams,, 2006) are a widely used choice for constructing emulators due to their flexibility, accuracy, and strong theoretical foundation. They are particularly well-suited for approximating the complex, non-linear relationships commonly encountered in real-world problems. A key advantage of GPs lies in their ability not only to provide mean predictions but also to quantify uncertainty through analytical expressions of posteriors. Moreover, GPs naturally interpolate training data, ensuring exact predictions at observed points. This interpolation property is particularly important for dynamic simulators, which are inherently deterministic, as it preserves the consistency of the emulator with the original simulator outputs.
Several GP emulators for dynamic simulators have been developed, ranging from classic methods to recent advancements (Girard et al.,, 2002; Bhattacharya,, 2007; Conti et al.,, 2009; Mohammadi et al.,, 2024), achieving significant performance improvements. For a detailed overview of dynamic simulator emulation, we refer readers to Mohammadi et al., (2024). Notably, Bhattacharya, (2007) proposed a one-step-ahead emulation approach, which assumes that the model output at time $t$ depends only on the output at the previous time point $t-1$, a property commonly referred to as the Markov property. Building on this, Mohammadi et al., (2019) enhanced the approach by incorporating input uncertainty and emulating the numerical flow map. Despite these advancements, many emulators still require intensive computations, such as Monte Carlo (MC) approximations (Conti et al.,, 2009; Mohammadi et al.,, 2019) or grid-based methods (Bhattacharya,, 2007).
@@
In this article, we propose a fast and more accurate emulation method for dynamic simulators by treating the emulator as self-coupled, adopting the one-step-ahead approach while overcoming the aforementioned limitations of MC approximations. The proposed method builds on the framework of Mohammadi et al., (2019) but replaces the approximated distribution with the exact mean and variance. These analytically tractable predictive posterior mean and variance, enabled by Linked GPs (Kyzyurova et al.,, 2018; Ming and Guillas,, 2021), facilitate the computation of exact posteriors, thus eliminating the need for time-consuming MC approximations. Linked GPs, originally developed for coupled computer models, allow for the analytical derivation of the mean and variance of GPs at random inputs. These closed-form expressions are achievable with commonly used kernels, such as the squared exponential kernel or the Matérn kernel (Stein,, 1999), with smoothness parameters $\nu=1.5$ and $\nu=2.5$. By leveraging this framework, the proposed model achieves computational efficiency without any loss of information, as the computational burdens of most existing methods primarily arise from the approximation step.
@@ 2.1 Gaussian processes
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ represent the computer code output as a deterministic black-box function of the input $\mathbf{x}$, where $\mathbf{x}\in\mathcal{X}\subseteq\mathbb{R}^{d}$, and the corresponding output is $y=f(\mathbf{x})$. Typically, a GP prior assumes that the simulation outputs $\mathbf{y}$ follow a multivariate normal distribution:
$\mathbf{y}\sim\mathcal{N}\left(\alpha(\mathbf{x}),\tau^{2}K(\mathbf{x},\mathbf{x}^{\prime})\right),$
where $\alpha(\mathbf{x})$ is the mean function, $\tau^{2}$ is a scale parameter, and $K(\mathbf{x},\mathbf{x}^{\prime})$ is a positive-definite kernel matrix. Popular choices for the mean function $\alpha(\mathbf{x})$ include linear $(1,\mathbf{x}^{T})\beta$, constant $\alpha$, or zero, where $\beta$ is a vector of $d+1$ regression coefficients. For the kernel $K$, the squared exponential kernel or Matérn kernel is commonly adopted. In this paper, we focus on the anisotropic squared exponential kernel $K(\mathbf{x},\mathbf{x}^{\prime})=\exp{\left(-\sum^{d}_{k=1}\frac{\left(x_{k}-x_{k}^{\prime}\right)^{2}}{\theta_{k}}\right)}$, where $(\theta_{1},\ldots,\theta_{d})$ are the lengthscale hyperparameters that represent the rate of decay of correlation in each input dimension.
The posterior prediction of GPs at a new input point $\mathbf{x}$, given the input design $X=(\mathbf{x}_{1},\ldots,\mathbf{x}_{n})\subset\mathbb{R}^{n\times d}$ and corresponding outputs $\mathbf{y}$, is given as:
$\mu(\mathbf{x})$ $=\alpha(\mathbf{x})+\mathbf{k}^{T}(\mathbf{x})\mathbf{K}^{-1}\left(\mathbf{y}-\alpha(X)\right),\quad\text{and}$
$\sigma^{2}(\mathbf{x})$ $=\tau^{2}\left(1-\mathbf{k}^{T}(\mathbf{x})\mathbf{K}^{-1}\mathbf{k}(\mathbf{x})\right),$
where $\mathbf{k}(\mathbf{x})_{i}=K(X,\mathbf{x})_{i}$ and $\mathbf{K}_{ij}=K(\mathbf{x}_{i},\mathbf{x}_{j})$. For more details about GPs in the context of computer experiments, we refer to Santner et al., (2018) and Gramacy, (2020).
The parameters $\left\{\beta,\tau^{2},\theta_{1},\ldots,\theta_{d}\right\}$ in the GP model can be estimated by maximizing the log-likelihood:
$-\frac{1}{2}\left(n\log 2\pi+n\log\tau^{2}+\log|\mathbf{K}|+\frac{1}{\tau^{2}}\left(\mathbf{y}-\alpha(X)\right)^{T}\mathbf{K}^{-1}\left(\mathbf{y}-\alpha(X)\right)\right).$
@@
$\hat{\beta}=\left(X_{n}^{T}\mathbf{K}^{-1}X_{n}\right)^{-1}X_{n}^{T}\mathbf{K}^{-1}\mathbf{y}\quad\text{and}\quad\hat{\tau}^{2}=\frac{\left(\mathbf{y}-(1,X_{n}^{T})\hat{\beta}\right)^{T}\mathbf{K}^{-1}\left(\mathbf{y}-(1,X_{n}^{T})\hat{\beta}\right)}{n}.$
In practice, optimization algorithms such as L-BFGS-B (Byrd et al.,, 1995; Zhu et al.,, 1997) are commonly employed to maximize the log-likelihood.
@@ 2.2 Linked GPs
Linked GP emulators (Kyzyurova et al.,, 2018; Ming and Guillas,, 2021) are designed to emulate two or more coupled systems of computer models. Consider $n_{1}$ computer models at the first layer, $f_{l}:\mathbb{R}^{d_{1}}\rightarrow\mathbb{R}$ for $l=1,\ldots,n_{1}$, which are connected to a second-layer model $g:\mathbb{R}^{d_{2}}\rightarrow\mathbb{R}$. In this framework, the outputs of $f_{1},\ldots,f_{n_{1}}$ serve as part of the input to $g$. Assuming that $f_{1},\ldots,f_{n_{1}}$ are independent and share the same set of inputs, the posterior predictive distribution at a new input $\mathbf{x}$, given the input-output pairs $\left(X_{n},f_{l}(X_{n})\right)_{l=1,\ldots,n_{1}}$ and $\left(Z_{n},g(Z_{n})\right)$ is expressed as:
$p$ $(g\circ(f_{1},\ldots,f_{n_{1}})(\mathbf{x})|\left\{f_{l}(X_{n})\right\}_{l=1,\ldots,n_{1}},g(Z_{n}),\mathbf{x})$
$=\int p(g\left(\mathbf{x}_{g},f_{1}(\mathbf{x}_{f}),\ldots,f_{n_{1}}(\mathbf{x}_{f})\right)|\left\{f_{l}(X_{n})\right\}_{l=1,\ldots,n_{1}},g(Z_{n}),\mathbf{x})$
×$\times\prod^{n_{1}}_{l=1}p(f_{l}(\mathbf{x}_{f})|f_{l}(X_{l}))df_{1}(\mathbf{x}_{f})\ldots df_{n_{1}}(\mathbf{x}_{f})$ (1)
where $\mathbf{x}_{g}$ denotes the unlinked inputs of $g$, and $\mathbf{x}_{f}$ represents the inputs for $f_{1},\ldots,f_{n_{1}}$, which are a subset of $\mathbf{x}$. This posterior is known to be analytically intractable (Girard et al.,, 2002). Numerical techniques, such as Markov Chain Monte Carlo, can approximate the posterior but are computationally demanding. To address this limitation, subsequent works replaced $f_{1},\ldots,f_{n_{1}}$ and $g$ with their respective GP emulators $\hat{f}_{1},\ldots,\hat{f}_{n_{1}}$ and $\hat{g}$ and applied Gaussian approximations. Using this approach, Kyzyurova et al., (2018) derived the closed-form expressions under the squared exponential kernel, and Ming and Guillas, (2021) extended the posterior to the Matérn kernels.
Linked GP emulators minimize the Kullback-Leibler divergence between the emulator and a Gaussian density (Ming and Guillas,, 2021). Moreover, they demonstrate superior performance compared to composite emulator, which disregard the coupled relationships and consider only the inputs of the first layer and the outputs of the final layer. This concept of linked GPs can be naturally extended to systems with multiple layers involving iterative or parallel structures (Ming and Guillas,, 2021; Heo and Sung,, 2023), high-dimensional output (Dolski et al.,, 2024), as well as deep GPs (Ming et al.,, 2023).
@@ 2.3 One-step-ahead emulations
Consider a dynamic computer simulator without forcing input $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times T}$ that produces $T$ outputs evolving over time. The simulator $f$ is governed by $d$ ODEs, where the state variables at time $t_{s}$ are expressed as $\mathbf{x}(t_{s})=(x_{1}(t_{s}),\ldots,x_{d}(t_{s}))^{T}$, for $s=0,\ldots,T$. The one-step-ahead approach emulates the system’s flow map over short time intervals, assuming the Markov property, which states that the state $\mathbf{x}(t_{s})$ depends only on the previous state $\mathbf{x}(t_{s-1})$. Once an emulator $\hat{f}$ is constructed based on the input-output pairs at the initial condition $\mathbf{x}(t_{0})$ and the corresponding output $\mathbf{x}(t_{1})$, predictions for future states $\hat{\mathbf{x}}(t_{s})$ for $s=1,\ldots,T$ can be obtained iteratively by leveraging the Markov assumption (Mohammadi et al.,, 2019, 2024):
$\hat{x}_{m}(t_{s+1})=\hat{f}_{m}(\mathbf{x}(t_{s})),$ (2)
@@
where $m=1,\ldots,d$. This approach is particularly effective for periodic systems, as it can accurately predict the next state based on the current state within the input range, making it ideal for capturing the system’s repetitive dynamics. Furthermore, the one-step-ahead approach provides an efficient, reliable, and computationally inexpensive alternative to other emulation methods for dynamic simulators (Conti et al.,, 2009; Stolfi and Castiglione,, 2021).
@@
Mohammadi et al., (2019) incorporate input uncertainty at each time step into the emulators by adopting MC methods and leveraging the laws of total expectation and total variance to approximate the posterior in (2.2):
$\mathbb{E}\left[\hat{x}_{m}(t_{s})\right]$ $=\mathbb{E}\left[\hat{f}_{m}(\mathbf{x}(t_{s-1}))\right]=\mathbb{E}\left[\mathbb{E}\left[\hat{f}_{m}(\mathbf{x}(t_{s-1}))|\mathbf{x}(t_{s-1})\right]\right],$ (3)
$\mathbb{V}\left[\hat{x}_{m}(t_{s})\right]$ $=\mathbb{V}\left[\hat{f}_{m}(\mathbf{x}(t_{s-1}))\right]$
$=\mathbb{E}\left[\mathbb{V}\left[\hat{f}_{m}(\mathbf{x}(t_{s-1}))|\mathbf{x}(t_{s-1})\right]\right]+\mathbb{V}\left[\mathbb{E}\left[\hat{f}_{m}(\mathbf{x}(t_{s-1}))|\mathbf{x}(t_{s-1})\right]\right].$ (4)
Modeling GPs for $f_{l}$ This framework enables the propagation of uncertainties inherent in the inputs through the emulator. Furthermore, the approach has been extended to account for correlated emulators, enabling a comprehensive treatment of dependencies across time steps.

### v2

@@ Abstract
Dynamic simulators are computational models governed by differential equations that evolve over time. They are essential for scientific and engineering applications but remain challenging to emulate because of the unpredictable behavior of complex systems. To address this challenge, this paper introduces a fast and accurate Gaussian Process (GP)-based emulation method for complex dynamic simulators. By integrating linked GPs into the one-step-ahead emulation framework, the proposed algorithm provides exact and tractable computation of the posterior mean and variance, solving a problem previously considered computationally intractable and eliminating the need for expensive Monte Carlo approximations. This approach substantially reduces computation time while maintaining or improving predictive accuracy. Furthermore, the method naturally extends to systems with forcing inputs by incorporating them as additional variables within the GP framework. Numerical experiments on multiple dynamic systems demonstrate the efficiency and computational advantages of the proposed approach. An R package, dynemu, which implements the one-step-ahead emulation approach, is available on CRAN.
@@
Junoh Heo***Corresponding author. These authors gratefully acknowledge funding from NSF DMS 2338018.
@@ 1 Introduction
Gaussian processes (GPs) (Rasmussen and Williams,, 2006) are a widely used choice for constructing emulators due to their flexibility and strong theoretical foundation. They are well-suited for approximating complex, nonlinear relationships commonly encountered in real-world problems, although their performance can be sensitive to the choice of kernel and modeling assumptions. A key advantage of GPs lies in their ability not only to provide mean predictions but also to quantify uncertainty through analytical posterior expressions. In deterministic settings, GPs interpolate training data, a property particularly important for dynamic systems governed by deterministic equations, as it ensures consistency between the emulator and the original simulator outputs.
Several GP emulators for dynamic simulators have been developed, ranging from classic methods to recent advancements (Girard et al.,, 2002; Bhattacharya,, 2007; Conti et al.,, 2009; Mohammadi et al.,, 2024). For a detailed overview of dynamic simulator emulation, we refer readers to Mohammadi et al., (2024). Notably, Bhattacharya, (2007) proposed a one-step-ahead emulation approach, which assumes that the model output at time $t$ depends only on the output at the previous time point t1t-1, a property commonly referred to as the Markov property. Building on this, Mohammadi et al., (2019) enhanced the approach by incorporating input uncertainty and emulating the numerical flow map. Despite these advancements, many emulators still require intensive computations, such as Monte Carlo (MC) approximations (Conti et al.,, 2009; Mohammadi et al.,, 2019) or grid-based methods (Bhattacharya,, 2007).
@@
In this article, we propose a fast and accurate emulation method for dynamic simulators by treating the emulator as self-coupled and adopting a one-step-ahead approach, thereby addressing the limitations of MC approximations. The proposed method builds on the framework of Mohammadi et al., (2019) but replaces the MC-based estimation of the posterior with exact closed-form expressions for the predictive mean and variance. These analytically tractable expressions are enabled by linked GPs (Kyzyurova et al.,, 2018; Ming and Guillas,, 2021), originally developed for coupled computer models. In the context of dynamic emulation, recursive prediction introduces uncertainty at each step, making the inputs to the emulator uncertain inputs (i.e., stochastic inputs represented as random variables) rather than fixed values. The linked GP formulation allows this uncertainty to be propagated analytically over time, eliminating the need for repeated sampling and significantly reducing computational cost. By leveraging this framework, our method achieves computational efficiency without compromising accuracy, as most of the computational burdens in existing methods arise from the approximation step.
@@ 2.1 Gaussian processes
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ denote a deterministic black-box function representing the output of a computer model. The input design matrix is given by $X=(\mathbf{x}_{1},\ldots,\mathbf{x}_{n})^{\top}\in\mathbb{R}^{n\times d}$, where each $\mathbf{x}_{i}\in\mathbb{R}^{d}$ is a design point. The corresponding outputs are collected in the vector $\mathbf{y}=(y_{1},\ldots,y_{n})^{\top}$ with each response defined by $y_{i}=f(\mathbf{x}_{i})$. Typically, a GP prior assumes that the simulation outputs $\mathbf{y}$ follow a multivariate normal distribution:
$\mathbf{y}\sim\mathcal{N}_{n}\left(\alpha(\cdot),\tau^{2}K(\cdot,\cdot)\right),$
where $\alpha(\cdot)$ is the mean function, $\tau^{2}$ is a scale parameter, and $K(\cdot,\cdot^{\prime})$ is a positive-definite kernel function. Popular choices for the mean function $\alpha(\mathbf{x})$ include linear $(1,\mathbf{x}^{\top})\beta$, constant $\alpha$, or zero, where $\beta$ is a vector of $d+1$ regression coefficients. In this paper, we adopt the linear mean function $\alpha(\mathbf{x})=(1,\mathbf{x}^{\top})\beta$. For the kernel $K$, the squared exponential kernel or Matérn kernel (Stein,, 1999) is commonly adopted. In this paper, we focus on the anisotropic squared exponential kernel $K(\mathbf{x},\mathbf{x}^{\prime})=\exp{\left(-\sum^{d}_{k=1}\frac{\left(x_{k}-x_{k}^{\prime}\right)^{2}}{\theta_{k}}\right)}$, where $\mathbf{x}=(x_{1},\ldots,x_{d})^{\top}$ and $\mathbf{x}^{\prime}=(x_{1}^{\prime},\ldots,x_{d}^{\prime})^{\top}$ are input vectors of size $d\times 1$, and $(\theta_{1},\ldots,\theta_{d})$ are the lengthscale hyperparameters representing the rate of correlation decay in each input dimension.
The posterior prediction of GPs at a new input point $\mathbf{x}$, given the input design $X$ and corresponding outputs $\mathbf{y}$, is given as:
$\mu(\mathbf{x})$ $=\alpha(\mathbf{x})+\mathbf{k}^{\top}(\mathbf{x})\mathbf{K}^{-1}\left(\mathbf{y}-\alpha(X)\right),\quad\text{and}$
$\sigma^{2}(\mathbf{x})$ $=\tau^{2}\left(1-\mathbf{k}^{\top}(\mathbf{x})\mathbf{K}^{-1}\mathbf{k}(\mathbf{x})\right),$
where $\mathbf{k}(\mathbf{x})_{i}=K(X_{i},\mathbf{x})$ and $\mathbf{K}_{ij}=K(\mathbf{x}_{i},\mathbf{x}_{j})$. For more details about GPs in the context of computer experiments, we refer to Santner et al., (2018) and Gramacy, (2020).
The model parameters $\left\{\beta,\tau^{2},\theta_{1},\ldots,\theta_{d}\right\}$ are estimated by maximizing the log-likelihood:
$-\frac{1}{2}\left(n\log 2\pi+n\log\tau^{2}+\log|\mathbf{K}|+\frac{1}{\tau^{2}}\left(\mathbf{y}-\alpha(X)\right)^{\top}\mathbf{K}^{-1}\left(\mathbf{y}-\alpha(X)\right)\right).$
@@
$\hat{\beta}=\left(X_{n}^{\top}\mathbf{K}^{-1}X_{n}\right)^{-1}X_{n}^{\top}\mathbf{K}^{-1}\mathbf{y}\quad\text{and}\quad\hat{\tau}^{2}=\frac{\left(\mathbf{y}-(1,X_{n}^{\top})\hat{\beta}\right)^{\top}\mathbf{K}^{-1}\left(\mathbf{y}-(1,X_{n})\hat{\beta}\right)}{n}.$
All hyperparameters in this paper are estimated by maximum likelihood using the L-BFGS-B algorithm (Byrd et al.,, 1995; Zhu et al.,, 1997).
@@ 2.2 Linked GPs
Linked GP emulators (Kyzyurova et al.,, 2018; Ming and Guillas,, 2021) are designed to emulate two or more coupled systems of computer models. Consider $n$ computer models at the first layer, $f_{l}$ for $l=1,\ldots,n$, which are connected to a second-layer model $g$. In this framework, the outputs from $f_{1},\ldots,f_{n}$ serve as part of the input to $g$. Assuming $f_{1},\ldots,f_{n}$ are independent, the posterior predictive distribution at a new input $\mathbf{x}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{n},\mathbf{x}_{g})$ where $\mathbf{x}_{l}$ is the input to $f_{l}$ and $\mathbf{x}_{g}$ is the unlinked input to $g$, is expressed as:
$p$ $(g\circ(f_{1},\ldots,f_{n})(\mathbf{x}))=\int p(g\left(\mathbf{x}_{g},f_{1}(\mathbf{x}_{1}),\ldots,f_{n}(\mathbf{x}_{n})\right))\prod^{n}_{l=1}p(f_{l}(\mathbf{x}_{l}))df_{1}(\mathbf{x}_{f})\ldots df_{n}(\mathbf{x}_{f}).$ (1)
This posterior is known to be analytically intractable (Girard et al.,, 2002). Numerical techniques, such as Markov Chain Monte Carlo, can approximate the posterior but are computationally demanding. To address this limitation, subsequent works replaced $f_{1},\ldots,f_{n}$ and $g$ with their respective GP emulators $\hat{f}_{1},\ldots,\hat{f}_{n}$ and $\hat{g}$ and applied Gaussian approximations. Using this approach, Kyzyurova et al., (2018) derived the closed-form expressions for the posterior mean and variance under the squared exponential kernel, while Ming and Guillas, (2021) extended these results to Matérn kernels.
Linked GP emulators minimize the Kullback-Leibler divergence between the emulator and a Gaussian density (Ming and Guillas,, 2021). Moreover, they demonstrate superior performance compared to the composite emulator, which disregards the coupled relationships and considers only the inputs of the first layer and the outputs of the final layer. This concept of linked GPs can be naturally extended to systems with multiple layers involving iterative or parallel structures (Ming and Guillas,, 2021; Heo and Sung,, 2023), high-dimensional output (Dolski et al.,, 2024), as well as deep GPs (Ming et al.,, 2023).
@@ 2.3 One-step-ahead emulations
Consider a dynamic computer simulator $f$ governed by a set of $d$ ODEs, where the state at time $t_{s}$ is denoted by $\mathbf{x}(t_{s})=(x_{1}(t_{s}),\ldots,x_{d}(t_{s}))^{\top}\in\mathbb{R}^{d}$, for $s=0,\ldots,T$. The simulator may also include $p$-dimensional forcing inputs $\mathbf{w}(t_{s})=(w_{1}(t_{s}),\ldots,w_{p}(t_{s}))^{\top}$, which are typically assumed to be known. The simulator maps the current state and forcing inputs to the next state,
$f(\mathbf{x}(t_{s}),\mathbf{w}(t_{s}))=\begin{bmatrix}f_{1}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s}))\\ \vdots\\ f_{d}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s}))\end{bmatrix}=\begin{bmatrix}x_{1}(t_{s+1})\\ \vdots\\ x_{d}(t_{s+1})\end{bmatrix}=\mathbf{x}(t_{s+1}),$
thereby generating $d$ output trajectories over $T$ time steps.
The one-step-ahead approach emulates the system’s flow map over short time intervals, assuming the Markov property: the next state $\mathbf{x}(t_{s+1})$ depends only on the current state $\mathbf{x}(t_{s})$ and the forcing input $\mathbf{w}(t_{s})$. Once a set of emulators $\{\hat{f}_{m}\}_{m=1}^{d}$ is constructed using training data $\{(\mathbf{x}^{i}(t_{0}),\mathbf{w}^{i}(t_{0})),\mathbf{x}^{i}(t_{1})\}_{i=1}^{n}$, future states $\mathbf{x}(t_{s+1})$ can be predicted recursively by leveraging the Markov assumption (Mohammadi et al.,, 2019, 2024). Importantly, this one-step-ahead approach requires only pairs of states one step apart rather than full trajectories, thereby substantially reducing the need to solve ODEs across all time steps. For notational simplicity, we write $\mathbf{x}(t_{s})$ for predicted states for $s\geq 1$. Under this convention, the recursive prediction is given by:
$x_{m}(t_{s+1})=\hat{f}_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s})),$ (2)
@@
for $m=1,\ldots,d$, and $s=0,\ldots,T-1$. This approach is particularly effective for periodic systems, as it can accurately predict the next state based on the current state within the input range, making it well-suited for capturing repetitive dynamics. Furthermore, the one-step-ahead approach provides an efficient, reliable, and computationally inexpensive alternative to other emulation methods for dynamic simulators (Conti et al.,, 2009; Stolfi and Castiglione,, 2021).
@@
Mohammadi et al., (2019) extends this framework by incorporating input uncertainty at each time step by adopting MC methods, and approximating the posterior in (1) via the laws of total expectation and total variance:
$\mathbb{E}\left[x_{m}(t_{s+1})\right]$ $=\mathbb{E}\left[\hat{f}_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s}))\right]=\mathbb{E}_{\mathbf{x}(t_{s})}\left[\mathbb{E}\left[\hat{f}_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s}))|\mathbf{x}(t_{s}),\mathbf{w}(t_{s})\right]\right],$ (3)
$\mathbb{V}\left[x_{m}(t_{s+1})\right]$ $=\mathbb{V}\left[\hat{f}_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s}))\right]$
$=\mathbb{E}_{\mathbf{x}(t_{s})}\left[\mathbb{V}\left[\hat{f}_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s}))|\mathbf{x}(t_{s}),\mathbf{w}(t_{s})\right]\right]+\mathbb{V}_{\mathbf{x}(t_{s})}\left[\mathbb{E}\left[\hat{f}_{m}(\mathbf{x}(t_{s}),\mathbf{w}(t_{s}))|\mathbf{x}(t_{s}),\mathbf{w}(t_{s})\right]\right].$ (4)
In practice, equations (3) and (2.3) are approximated by drawing $n_{\mathrm{MC}}$ random samples of the uncertain inputs at every prediction step and evaluating the GP emulator at each sample. This procedure is repeated sequentially across all $T$ prediction steps, causing the total computational cost to increase with $n_{\mathrm{MC}}$ and $T$. While this framework effectively propagates input uncertainties through the emulator and has been extended to account for correlated emulators, Mohammadi et al., (2019) considers only systems without forcing inputs, limiting its flexibility for simulators with external covariates.